OpenStudy (anonymous):

the kinetic energy of a moving object varies directly with the square of its velocity. if a car is moving at 90 mph has kinetic energy 100 units, how many units of kinetic energy does the same car have at 10 mph?

OpenStudy (anonymous):

KE = k*v^2 k = KE/v^2 = 100/810 KE | v = 10 = 100/810 * 100 = 10000/810 units

OpenStudy (anonymous):

so its 810?

OpenStudy (anonymous):

or 10000/810?

OpenStudy (anonymous):

10000/810

OpenStudy (anonymous):

great thank you so much